# Emergent photons and mechanisms of confinement

### v1

@@ I Introduction
Yet both of these theories and regimes are linked to an exotic symmetry called the 1-form symmetry that has been of particular interest in the recent decade, mostly starting with the works [1, 2, 3] [^1] The 1-form symmetries were known in $\text{SU}(N)$ Yang-Mills theory at least since the late 70s [4], but until recently were relatively poorly understood.. On one hand, in QCD, quark confinement is associated with the would-be $Z_{N}$ 1-form symmetry — a symmetry of the pure glue sector — while the massless photons of QED are seen as Goldstone bosons of a $U(1)$ 1-form symmetry. [2, 5]. Therefore, both the massless photon regime and the nuclear confined regime can potentially be understood as phases of matter within the Ginzburg-Landau paradigm for 1-form symmetries (see also [6]).
In [7], it was argued that theories with $\mathbb{Z}_{N}$ 1-form symmetry in 3+1D have a generic photon phase for any $N$. This phase is seen as an enhancement of the 1-form symmetry $\mathbb{Z}_{N}\rightarrow\text{U}(1)$, and consequent spontaneous breaking of it, leading to a massless Goldstone boson — the photon.
@@
In parallel, the 3+1D $\mathbb{Z}_{N}$ LGT is a toy model of pure $SU(N)$ Yang-Mills theory, having both confined and deconfined phases. This model offers a deceptively simple picture of confinement as originating from the proliferation of string-like objects analogous to center vortices of $\text{SU}(N)$ Yang-Mills theory (see [10, 11] for a review). However, the existence of the photon phase requires this picture to be revised by also considering monopole particles [12, 7].
It is worth noting that the occurrence of a robust massless phase without a continuous global symmetry is surprising from the perspective of ordinary 0-form symmetries, as there can be no Goldstone bosons [^2] Alternatively, one could hit a conformal fixed point, but this typically requires fine-tuning.. Moreover, continuous symmetry cannot simply emerge, as any explicit breaking of an emergent 0-form continuous symmetry will lift the mass of the Goldstone boson.
That the same conclusion does not hold for $\mathbb{Z}_{N}$ 1-form symmetry in 3+1D is what makes the Millennium Problem of whether $\text{SU}(N)$ Yang-Mills theory has a mass gap [13] nontrivial. Indeed, based on the $\mathbb{Z}_{N}$ 1-form global symmetry group, we may be erroneously tempted to conclude that a gapless phase is untenable. Incidentally, in 2+1D $\text{SU}(N)$ Yang-Mills theory, the photon phase is generically unstable [7], resulting in the necessity of a mass gap under all generic symmetry-preserving deformations (see e.g. [14, 15, 16, 17]).
@@ II The $\mathbb{Z}_{N}$ Lattice Gauge Theory (LGT)
The model we consider here is a $Z_{N}$ Lattice Gauge Theory, which is to $Z_{N}$ 1-form symmetry what the Ising model is to an ordinary $Z_{2}$ symmetry. Like the Ising model, it is a theory with discrete $N$-state spins $s_{\ell}=0,1,2,\cdots,N-1$ on lattice links $\ell$. The spins $s_{\ell}$ naturally form a $\mathbb{Z}_{N}$ phase $U_{\ell}=e^{\frac{2\pi is_{\ell}}{N}}$. An action that dictates the dynamics of these spins can be constructed by defining $W_{p}=\prod_{\ell\in\partial p}U_{\ell}$ — a plaquette sized Wilson loop — where $\partial p$ is the set of boundary links of a plaquette $p$, inheriting the orientation of the plaquette (with the convention $U_{\ell}=U_{-\ell}^{*}$). Explicitly, consider a hyper-cubic lattice, and let $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$ be the boundary links of $p$ as in Fig. 1, then $W_{p}=U_{\ell_{1}}U_{\ell_{2}}U_{\ell_{3}}U_{\ell_{4}}$. The standard Wilson action and the partition function are given by
@@
A crucial property of this action in general dimension $d$ is that it is invariant under the $\mathbb{Z}_{N}$ 1-form symmetry [2]. To understand this symmetry applied to (1), consider an oriented $(d-1)$-dimensional hypersurface $H$ constructed out of $(d-1)$-cells of the dual lattice (an example for $d=2$ is given in Fig. 1). This will intersect a set of lattice links, which we collect into a set $L$ (marked blue in Fig. 1) and orient compatibly relative to $H$. A $\mathbb{Z}_{N}$ 1-form symmetry is generated by shifting [^3] We note that by convention $s_{-\ell}=-s_{\ell}\bmod N$. $s_{\ell}\rightarrow s_{\ell}+1\bmod N$ whenever $\ell\in L$; this is clearly an invariance of the action.
By taking $H$ to be such that it surrounds a single lattice site, the resulting transformation is the conventional gauge transformation.
Any contractible $H$ thus corresponds to a composition of local gauge transformations. On the other hand, non-contractible choices of $H$ that differ by a pure boundary yield the same transformation up to gauge symmetry. Such choices of $H$, up to pure boundary terms, are what is meant by a 1-form symmetry, so the global symmetry in question takes values in $H_{d-1}(\mathbb{Z}_{N})\cong H^{1}(\mathbb{Z}_{N})$.
The operators charged under the 1-form symmetry are Wilson loops $W[\mathcal{C}]=\prod_{\ell\in\mathcal{C}}U_{\ell}$ where $\mathcal{C}$ is a contour made out of the links of the lattice. It is easy to see that under the $1$-form symmetry transformation of Fig. 1b) the Wilson loop transforms as $W[\mathcal{C}]\rightarrow W[\mathcal{C}]\exp\left({i\frac{2\pi I(H,\mathcal{C})}{N}}\right)$ where $I(H,\mathcal{C})$ is the net intersection number between $H$ and $\mathcal{C}$. The fact that in the continuum the charged operator is a Wilson loop, which is an integral over the 1-form gauge field $A_{\mu}$ is what gives the 1-form symmetry its name.
The naive Landau paradigm would then indicate at least two phases: the $\mathbb{Z}_{N}$ restored (confined) and the spontaneously (symmetry) broken (SSB) phase, which can be seen respectively in the limits $\beta\rightarrow 0,\infty$. When $0<\beta\ll 1$, the contractible Wilson loops $W[\mathcal{C}]$ get a leading contribution at order $\beta^{A}$, where $A$ is the area of a minimal surface $S$ whose boundary is $\mathcal{C}$. By cluster decomposition, this implies that the non-contractible Wilson loops must have zero expectation value, i.e., the $\mathbb{Z}_{N}$ 1-form symmetry is restored. In the limit $\beta\gg 1$, the dominant configurations obey $W_{p}=1$, indicating that non-contractible Wilson loops get a nonzero expectation value.
@@
Refer to caption Refer to captionRefer to captionRefer to caption
In fact, a picture of the photon and the confined phase was first briefly touched on in Appendix B of the little-known paper by Michael Stone [12] and recently in [7], which we aim to confirm here. In this framework, both the photon and confined phases involve proliferating center vortices, but differ in the behavior of monopole-junctions: these remain heavy in the photon phase, while they proliferate in the confined phase. In this work, we confirm this idea.
@@ III Numerical simulations and results
Refer to captionRefer to caption
@@
In this section, we numerically study the two different classes of theories: $N\geq 5$ and $52<N<$ 5. We choose $N=7$ to represent the $N\geq 5$ regime in which three phases are well established [9, 8]. The goal is to associate these phases with (non-)proliferating center vortices and monopole-junctions. To this end, we begin by formally defining these objects on the lattice.
$-\frac{N}{2}<f_{p}\leq\frac{N}{2}\;.$ (2)
In Fig. 1c), we give an example of the center vortex configuration by shading the plaquettes for which $f_{p}\equiv 1\bmod N$. If $(df)_{c}=0$ were to hold, then the center vortex would form closed string worldsheet configurations on the dual lattice. However because $(df)_{c}\equiv 0\mod N$, there can be configurations for which $(df)_{c}=N(dm)_{c}$, where $m_{p}\in\mathbb{Z}$, so $(dm)_{c}\neq 0$ (a loop on the dual lattice) is our definition of a monopole-junction — a worldline where $N$ center-vortex worldsheets meet.
Consider now the $\mathbb{Z}_{7}$ LGT with the usual Wilson action (1). This theory was simulated on the 4-torus that we interpret as the Euclidean space-time of size $L^{4}$. To probe this model, we ran Monte Carlo simulations for 200 values of $\beta\in[0.5,2.7]$, and performed $2\cdot 10^{4}$ measurements each, with 20 decorrelation sweeps between measurements. Furthermore, we chose $L=10$ and estimated the error via bootstrap and present the results in Fig. 2. Note that the errors are too small to be seen.
This theory has three phases, which we can distinguish by the Polyakov loop (i.e. the Wilson loop winding along one of the cycles of the torus). Fig. 2 shows the Polyakov loop expectation value, which shows a clear jump at two values of $\beta_{1}\approx 1.157\pm 0.006$ and $\beta_{2}\approx 2.108\pm 0.006$. The confined phase is clear from the plot, as the Polyakov loop expectation value is (nearly) zero, but the middle (the photon) and the large $\beta$ (the SSB) phases both have a finite Polyakov loop VEV. To distinguish the two, we measure the following. Let $x=(t,x_{1},x_{2},x_{3})$ parametrize the 4D position on a torus (i.e. $t\sim t+L$ and $x_{i}\sim x_{i}+L$). Let $\mathcal{L}_{\bm{x}}$ be the set of all links at the spatial point $\bm{x}=(x_{1},x_{2},x_{3})$ pointing in the temporal direction. Now, define the temporal Polyakov loop to be $P_{\bm{x}}=\prod_{\ell\in\mathcal{L}_{\bm{x}}}U_{\ell}$. For a given configuration, we define the spatially smeared Polyakov loop $\bar{P}=\frac{1}{L^{3}}\sum_{\bm{x}}P_{\bm{x}}$. The $\bar{P}$ is a complex number that depends on the configuration.
In the inset of Fig. 2, we plot the Monte Carlo histogram of this observable. One can clearly see that for the SSB phase, the smeared Wilson loop shows a $\mathbb{Z}_{7}$ symmetric distribution reflecting a spontaneously broken $\mathbb{Z}_{7}$ 1-form symmetry, while in the photon phase, a circular distribution emerges, indicating the emergence of the $\text{U}(1)$ 1-form symmetry. In the confined phase, the distribution of $\bar{P}$ is unsurprisingly sharply peaked around zero. In the inset of Fig. 2, we also show 3D slices of typical configurations in the three phases. The string-like objects are illustrations of the center vortices. One can clearly see that they form short loops (or rather, short worldsheets in 4D) in the $\beta>\beta_{2}$ SSB phase, while they clearly proliferate in both the photon and the confined phase. On the other hand, the monopole-junctions, shown as little cubes on 3D slices, are nonexistent in the SSB phase, are very rare (i.e. form short disconnected loops) in the photon phase, and are abundant in the confined phase. This establishes the connection between center vortices and monopole-junctions as proposed in [7].
This behavior is expected for any $N\geq 5$. However, in [22], a $\mathbb{Z}_{4}$ theory was considered, and there the authors considered a model with an additional parameter, $\tilde{\beta}$ which coupled two charge 2 Wilson loops, i.e. the action was modified to

### v2

@@ I Introduction
Yet both of these theories and regimes are linked to an exotic symmetry called the 1-form symmetry that has been of particular interest in the past decade, mostly starting with the works [1, 2, 3] [^1] The 1-form symmetries were known in $\text{SU}(N)$ Yang-Mills theory at least since the late 70s [4], but until recently were relatively poorly understood.. On one hand, in QCD, quark confinement is associated with the would-be $\mathbb{Z}_{N}$ 1-form symmetry — a symmetry of the pure glue sector — while the massless photons of QED are seen as Goldstone bosons of a $\text{U}(1)$ 1-form symmetry [2, 5]. Therefore, both the massless photon regime and the nuclear confined regime can potentially be understood as phases of matter within the Ginzburg-Landau paradigm for 1-form symmetries (see also [6]).
In [7], it was argued that theories with $\mathbb{Z}_{N}$ 1-form symmetry in 3+1D have a generic photon phase for any $N$. This phase is seen as an enhancement of the 1-form symmetry, $\mathbb{Z}_{N}\rightarrow\text{U}(1)$, and consequent spontaneous breaking of it, leading to a massless Goldstone boson: the photon.
@@
In parallel, the 3+1D $\mathbb{Z}_{N}$ LGT is a toy model of pure $\text{SU}(N)$ Yang-Mills theory, having both confined and deconfined phases. This model offers a deceptively simple picture of confinement as originating from the proliferation of string-like objects analogous to center vortices of $\text{SU}(N)$ Yang-Mills theory (see [10, 11] for a review). However, the existence of the photon phase requires this picture to be revised by also considering monopole particles [12, 7].
The occurrence of a robust exactly massless Goldstone phase without an exact continuous symmetry may be surprising from the point of view of 0-form symmetry intuition (see SM for a review). This robustness in 3+1D is what makes the Millennium Problem of whether $\text{SU}(N)$ Yang-Mills theory has a mass gap [13] nontrivial. Indeed, based on the $\mathbb{Z}_{N}$ 11-form global symmetry group, we may be erroneously tempted to conclude that a gapless phase is untenable. That a robust photon phase exists invalidates this lore.
Incidentally, in 2+1D the lore is true, as the photon phase is generically unstable [7], resulting in the necessity of a mass gap under all generic symmetry-preserving deformations (see e.g. [14, 15, 16, 17]).
@@ II The $\mathbb{Z}_{N}$ Lattice Gauge Theory (LGT)
The model we consider here is a $\mathbb{Z}_{N}$ Lattice Gauge Theory, which is to $\mathbb{Z}_{N}$ 1-form symmetry what the Ising model is to an ordinary $\mathbb{Z}_{2}$ symmetry. It is a theory with discrete $N$-state spins $s_{\ell}=0,1,2,\cdots,N-1$ on lattice links $\ell$. The spins $s_{\ell}$ naturally form a $\mathbb{Z}_{N}$ phase $U_{\ell}=e^{\frac{2\pi is_{\ell}}{N}}$. An action that dictates the dynamics of these spins can be constructed by defining $W_{p}=\prod_{\ell\in\partial p}U_{\ell}$ — a plaquette-sized Wilson loop — where $\partial p$ is the set of boundary links of a plaquette $p$, inheriting the orientation of the plaquette (with the convention $U_{\ell}=U_{-\ell}^{*}$). The standard Wilson action and the partition function are then given by
@@
A crucial property of this action in general dimension $d$ is that it is invariant under the $\mathbb{Z}_{N}$ 1-form symmetry [2]. To understand this symmetry applied to (1), consider an oriented (d1)(d-1)-dimensional hypersurface $H$ constructed out of (d1)(d-1)-cells of the dual lattice (an example for $d=2$ is given in Fig. 1b). This will intersect a set of lattice links, which we collect into a set $L$ (marked blue in Fig. 1b) and orient compatibly relative to $H$. A $\mathbb{Z}_{N}$ 1-form symmetry is generated by shifting [^2] We note that by convention $_{-\ell}=-s_{\ell}\bmod N. ss$$_{\ell}\rightarrow s_{\ell}+1\bmod N$ whenever $\ell\in L$; this is clearly an invariance of the action.
It is easily checked that for $H$ contractible, the resulting invariance of action is the conventional gauge transformations, while non-contractible choices of $H$ correspond to a global $1$-form symmetry. Thus, the transformations associated with homotopic choices of $H$ are gauge-equivalent, and so the $1$-form symmetry is valued in the co-dimension 1 homology of the space-time manifold with coefficients in $\mathbb{Z}_{N}$.
The operators charged under the 1-form symmetry are Wilson loops $W[\mathcal{C}]=\prod_{\ell\in\mathcal{C}}U_{\ell}$ where $\mathcal{C}$ is a contour made out of the links of the lattice. It is easy to see that under the $1$-form symmetry transformation for a hypersurface $H$ the Wilson loop transforms as $W[\mathcal{C}]\rightarrow W[\mathcal{C}]\exp\left({i\frac{2\pi I(H,\mathcal{C})}{N}}\right)$ where $I(H,\mathcal{C})$ is the net intersection number between $H$ and $\mathcal{C}$. The fact that in the continuum the charged operator is a Wilson loop, which is an integral over the 1-form gauge field $A_{\mu}$ is what gives the 1-form symmetry its name.
The naive Landau paradigm would then indicate at least two phases: the $\mathbb{Z}_{N}$ restored (confined) and the spontaneously symmetry broken (SSB) phase, which can be seen respectively in the limits $\beta\rightarrow 0,\infty$. When $0<\beta\ll 1$, the contractible Wilson loops $W[\mathcal{C}]$ get a leading contribution at order $\beta^{A}$, where $A$ is the area of a minimal surface $S$ whose boundary is $\mathcal{C}$. By cluster decomposition, this implies that the non-contractible Wilson loops (i.e. Polyakov loops) must have zero expectation value, i.e., the $\mathbb{Z}_{N}$ 1-form symmetry is restored. In the limit $\beta\gg 1$, the dominant configurations obey $W_{p}=1$, indicating that non-contractible Wilson loops get a nonzero expectation value.
@@
In fact, a picture of the photon and the confined phase was first briefly touched on in Appendix B of the little-known paper by Stone [12] and recently in [7], which we aim to confirm here. In this framework, both the photon and confined phases involve proliferating center vortices, but differ in the behavior of monopole-junctions: these remain heavy in the photon phase, while they proliferate in the confined phase.
@@ III Numerical simulations and results
Refer to captionRefer to captionRefer to captionRefer to caption
In this section, we numerically study the two different classes of theories: $N\geq 5$ and $2<N<5$. We choose $N=7$ to represent the $N\geq 5$ regime in which three phases are well established [9, 8]. The goal is to associate these phases with (non-)proliferating center vortices and monopole-junctions. To this end, we begin by formally defining these objects on the lattice.
$\frac{N}{2}<f_{p}\leq\frac{N}{2}\;.$ (2)
In Fig. 1c), we give an example of a center vortex configuration by shading the plaquettes for which $f_{p}\equiv 1\bmod N$. If $(df)_{c}=0$ were to hold, then the center vortex would form closed string worldsheet configurations on the dual lattice. However, because $(df)_{c}\equiv 0\mod N$, there can be configurations for which $(df)_{c}=N(dm)_{c}$, so $(dm)_{c}\neq 0$ (a loop on the dual lattice) is our definition of a monopole-junction – a worldline where $N$ center-vortex worldsheets meet.
Consider now the $\mathbb{Z}_{7}$ LGT with the usual Wilson action (1). To probe this model, we ran Monte Carlo simulations for 200 values of $\beta\in[0.5,2.7]$, and performed $2\cdot 10^{4}$ measurements each, with 20 decorrelation sweeps between measurements (see SM for more details on the simulations). Furthermore, we chose $L=10$ and estimated the error via bootstrap. We present the results in Fig. 2. Note that the errors are too small to be seen.
This theory has three phases, which we can distinguish by the Polyakov loop (i.e. the Wilson loop winding along one of the cycles of the torus). Fig. 2 shows the Polyakov loop expectation value, which shows a clear 1st order jump at two values of $\beta_{1}\approx 1.157\pm 0.006$ and $\beta_{2}\approx 2.108\pm 0.006$. The confined phase appears for small values where the Polyakov loop expectation value is (nearly) zero, but the middle (the photon) and the large $\beta$ (the SSB) phases both have a finite Polyakov loop VEV. To distinguish the two, we measure the following. Let $x=(t,x_{1},x_{2},x_{3})$ parametrize the 4D position on a torus (i.e. $t\sim t+L$ and $x_{i}\sim x_{i}+L$). Let $\mathcal{L}_{\bm{x}}$ be the sequence of all links pointing in the temporal direction at the spatial point $\bm{x}=(x_{1},x_{2},x_{3})$. Now, define the temporal Polyakov loop to be $P_{\bm{x}}=\prod_{\ell\in\mathcal{L}_{\bm{x}}}U_{\ell}$. To each configuration we assign a spatially smeared Polyakov loop $\bar{P}=\frac{1}{L^{3}}\sum_{\bm{x}}P_{\bm{x}}$, and plot its distribution across the Monte Carlo measurements for the typical representatives of the three phases in Fig. 2, as well as 3D slices of the corresponding configurations.
In the SSB phase $\beta>\beta_{2}$, the smeared Polyakov loop shows a $\mathbb{Z}_{7}$ symmetric distribution reflecting a spontaneously broken $\mathbb{Z}_{7}$ 1-form symmetry. In this phase, the center vortices correspond to short loops. As $\beta$ is lowered to $\beta_{1}\lesssim\beta\lesssim\beta_{2}$, the system transitions into the photon phase, the center vortices proliferate, and $\bar{P}$ develops a circular distribution, indicating the emergence of the $\text{U}(1)$ 1-form symmetry. Notably, the monopoles – depicted by green and yellow cubes in the figure – are uncommon, indicating that monopole worldlines are short and non-proliferating. As $\beta$ is reduced further to $0<\beta<\beta_{1}$, the confined phase sets in, whose hallmark is a distribution of $\bar{P}$ unsurprisingly sharply peaked around zero, and a proliferation of both center vortices as well as monopoles, consistent with the proposal of [7].
This behavior is expected for any $N\geq 5$. However, in [22, 23, 24], a $\mathbb{Z}_{4}$ theory was studied, with the following modified action
$S=-\beta\sum_{p}\text{Re\;}W_{p}-\tilde{\beta}\sum_{p}\text{Re\;}W_{p}^{2}\;.$ (3)
They argued that for a particular choice of $\beta$ and $\tilde{\beta}$, a massless photon phase emerges. Indeed in Fig. 3 we present the Monte Carlo time series histogram of the smeared Polyakov loops in Fig. 3. As can be observed, we see the same pattern of circle formation indicating a photon phase in between the confined and the SSB phase, confirming the photon phase.
Figure 3: Histogram of the smeared Polyakov loop as a function of $\beta$ and $\tilde{\beta}$ (see (3)) of the $\mathbb{Z}_{4}$ lattice gauge theory.
Now we turn to the $\mathbb{Z}_{3}$ model. In this model with the standard Wilson action (1), there are only two phases. The picture of this model is that as $\beta$ decreases from $\beta\gg 1$ to $\beta\ll 1$, both center vortices and monopole-junctions become abundant simultaneously. To open a photon phase, we therefore add a monopole suppression term to the action:
$S=-\beta\sum_{p}\text{Re\;}W_{p}+\mu\sum_{c}(dm)_{c}^{2}\;.$ (4)
To identify phase transitions, we look at the action susceptibility (“heat capacity”) defined as
In Fig. 4 we plot numerical Monte Carlo results for $\chi_{S}$ as a function of $\beta$ at $\mu=1$. We ran Monte Carlo simulations for $L=8,10,12,$ and $14$ for various values of $\beta$ (for simulation details see SM). The results for $\chi_{S}$ can be seen in Fig. 4. As is clear, the susceptibility shows two peaks, indicating a three-phase scenario, with phase transitions at $\beta_{1}\approx 0.265$ and $\beta_{2}\approx 0.512$. The inset shows the smeared Polyakov loop $\bar{P}$ for values of $\beta$ close to the second transition, showing the circles forming in the photon phase.
[Uncaptioned image][Uncaptioned image]
@@
Finally, we also consider the connected plaquette-plaquette correlators (PPC) defined as (no sum over $\mu,\nu$)
$C^{\pm}_{x-y;\mu\nu}=\left\langle W_{p_{x,\mu\nu}}^{\pm 1}W_{p_{y,\mu\nu}}\right\rangle-\big\langle W_{p_{x,\mu\nu}}^{\pm 1}\big\rangle\big\langle W_{p_{y,\mu\nu}}\big\rangle\;,$ (6)
where $W_{p}^{-1}=1/W_{p}$, and where $p_{x,\mu\nu}$ indicates a plaquette with a corner at lattice site $x$, expanding in directions $\mu,\nu=0,1,2,3$, and averages are over ensemble and all spatial translations. At long distances in the photon phase, the theory flows to a pure $\text{U}(1)$ gauge theory with the action $S=\frac{1}{4}\int d^{4}x\;F_{\mu\nu}F^{\mu\nu}$, which describes free photons. All gauge invariant local operators of the $\text{U}(1)$ gauge theory must be built from the field-strength operator $F_{\mu\nu}$, so we expect $C_{x-y;\mu\nu}\propto\left\langle F_{\mu\nu}(x)F_{\mu\nu}(y)\right\rangle$ for large |xy||x-y|. Further, we have that
where $D_{\mu\nu;\rho\sigma}$ is independent of the modulus |xy||x-y| and is easily computed [^3] One way is to note that $\left\langle A_{\mu}(x)A_{\nu}(y)\right\rangle=\delta_{\mu\nu}/|x-y|^{2}$ in the Feynman gauge, and then take derivatives., but is not relevant for our discussion. Let us define a combination [^4] The combination is chosen because, by reflection positivity, it is a sum of positive definite parts.
$C(n)=\!\!\sum_{\begin{subarray}{c}\rho,\mu,\nu\\ \mu>\nu\neq\rho\end{subarray}}C^{-}_{n\hat{\rho},\mu\nu}\!+\!\sum_{\begin{subarray}{c}\rho,\mu\\ \rho\neq\mu\end{subarray}}C^{+}_{n\hat{\rho},\rho\mu}\approx\underbrace{\frac{K}{n^{4}}\!+\!\frac{K}{(L-n)^{4}}}_{C_{\text{asym}}(n)}\;,$ (8)
where $\hat{\rho}$ is a lattice unit vector in the direction $\rho$, and $K$ is a constant, and where we took into account the corrections due to the finiteness of the periodic lattice.
Since in the ratio $C(n)/C(n+1)$ the constant $K$ cancels, all correlators in a Coulomb phase should asymptote to a universal function. In Fig. 5 we show numerical results of this ratio for $\mathbb{Z}_{3},\mathbb{Z}_{4}$ and $\mathbb{Z}_{7}$ in the Coulomb phase and find excellent asymptotic agreement, confirming the photon phase of the model. In the SM, we also show independent correlators $C^{\pm}(n)$ and find that they perfectly fit the expansion into powers of $1/n^{2}$.
Figure 5: The correlator ratio $C(n)/C(n+1)$ in the photon phases of $\mathbb{Z}_{3},\mathbb{Z}_{4}$, and $\mathbb{Z}_{7}$ on an $L=16$ lattice compared to the long-distance expectation (see (8)). Insets: Corresponding smeared Polyakov loops.
IV Conclusions
We numerically showed that $\mathbb{Z}_{N}$ lattice gauge theories have three phases for all $N\geq 3$, supplementing the spontaneously broken and restored $\mathbb{Z}_{N}$ 1-form symmetry with a phase of emergent photons where the symmetry enhances to $\text{U}(1)$.
In addition, we demonstrated a direct link between the photon phase and the proliferation of string-like center vortices, while the confined phase was linked to proliferation of both vortices and vortex junctions (monopoles) as proposed in [7]. This suggests that center vortices are not directly responsible for confinement, as commonly believed for $\text{SU}(N)$ gauge theory. [^5] Defining monopoles and center vortices requires gauge fixing in $\text{SU}(N)$ gauge theories (see e.g. [25, 10, 11]), although there are some efforts to identify such objects without it [26, 26]. Further, a novel proposal of how to discretize $\text{SU}(N)$ gauge theories [27, 28], may be related to a gauge invariant definition of vortices and monopoles [29].
The case of $\mathbb{Z}_{2}$ lattice gauge theory is more subtle as it does not have a global charge conjugation symmetry CC. [^6] TS would like to thank M. Nguyen and E. Richards for discussions on this point. While it was proposed in [7] to have a photon phase when deformed, such deformations would introduce the $C$-symmetry. A similar issue occurs in dihedral group ($D_{N}$) lattice gauge theories, which have no $C$ symmetry and only at most a $\mathbb{Z}_{2}$ 1-form symmetry. Yet for large enough $N$, $D_{N}$ LGT should be indistinguishable from $O(2)$ LGT, which certainly has a photon phase, albeit with a non-invertible emergent 1-form symmetry [^7] The lattice simulations of $D_{N}$ seem to indicate three phases for $N\gtrsim 6$ [30].. The Landau paradigm for $\mathbb{Z}_{2}$ symmetry therefore likely involves a massless phase as suggested in [7], although the details are more subtle and are left for future explorations.
Acknowledgement
We would like to thank Mathew Bullimore, Xavier Crean, Christof Gattringer, Tyler Helmuth, Nabil Iqbal, Zohar Komargodski, Biagio Lucini, Mendel Nguyen, Emily Richards, Shu-Heng Shao and Yuya Tanizaki for the discussions and comments. This work has made use of the Hamilton HPC Service of Durham University. TS is supported by the Royal Society University Research Fellowship and, in part, by the STFC consolidated grant number ST/T000708/1. JG is supported by EPSRC grant EP/Y028872/1.
Supplementary Material
Appendix A Details of simulations
All measurements were generated on a 4-torus of lattice size $L^{4}$, interpreted as Euclidean space-time, using Monte Carlo simulations with sequential sweeps of single-link Metropolis updates, where each link is updated once per sweep. A trial link value, chosen uniformly from $\mathbb{Z}_{N}$, is accepted with probability $\min\{1,e^{-\Delta S}\}$ based on the local action change $\Delta S$; otherwise, the original value is retained. Table 1 shows the details of lattice simulations presented in the Figures of the main text.
Figure 2 Figure 3 Figure 4 Figure 5
Theory $\mathbb{Z}_{7}$ $\mathbb{Z}_{4}$ $\mathbb{Z}_{3}$ $\mathbb{Z}_{3}$, $\mathbb{Z}_{4}$, $\mathbb{Z}_{7}$
Lattice sizes 1010 66 $8,10,12,14$ 1616
# measurements $2\times 10^{4}$ $10^{4}$ $4\times 10^{4}$ $1.28\times 10^{6}$
Decor. sweeps 20 1 100 500
Table 1: Simulation details.
Appendix B Binned Jackknife Error Estimation
Consecutive Monte Carlo samples are generally correlated, so straightforward error estimates tend to underestimate uncertainties. To address this, we divide the data into $N_{b}$ contiguous bins of size $B$ [31], chosen large enough such that correlations between the bins can be neglected. Jackknife samples are then formed by omitting one bin at a time and averaging over the remaining $N_{b}-1$ bins, with the variance of the resulting estimates rescaled by $(N_{b}-1)/N_{b}$ to compensate for the fact that each jackknife sample is based on $N_{b}-1$ bins rather than the full set, which would otherwise bias the variance downward. Let $\mathcal{O}$ be any observable and denote by $\overline{\mathcal{O}}_{B,1},\dots,\overline{\mathcal{O}}_{B,N_{b}}$ the bin averages. The jackknife estimate from leaving out bin $i$ and the jackknife mean are given by
$\mathcal{O}_{i}=\frac{1}{N_{b}-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N_{b}}\overline{\mathcal{O}}_{N,j}\ \ \text{ and }\ \ \overline{\mathcal{O}}=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathcal{O}_{i}.$
The statistical error can then be obtained by:
$\sigma_{\mathcal{O}}^{2}=\frac{N_{b}-1}{N_{b}}\sum_{i=1}^{N_{b}}(\mathcal{O}_{i}-\bar{\mathcal{O}})^{2},$
In practice, the bin size is increased successively, $B=2,5,10,20,50,100,200,500,\dots$, and once the error saturates the corresponding binned averages can be treated as effectively independent.
This saturation point also provides a useful heuristic for approximating how many sweeps should be skipped between recorded configurations in order to obtain uncorrelated samples. Effectively, we run the simulation with a given observable in mind, monitor the jackknife error as the bin size increases, and use the onset of a plateau to assess autocorrelation. If the plateau only emerges at large bin sizes, this indicates that configurations remain correlated over many sweeps, and the number of skipped sweeps should be increased accordingly. By doing so, we avoid recording correlated configurations, thereby reducing memory and computational costs while maintaining statistical accuracy. The evolution of the binning error for the longitudinal correlators can be seen in Figure 6.
Figure 6: Saturating binning error for longitudinal correlators. The dashed line marks the onset of the plateau, motivating a choice of 500 decorrelation sweeps.
Appendix C Smeared Polyakov Loops
Recall that the temporal Polyakov loop was defined to be $P_{\bm{x}}=\prod_{l\in\mathcal{L}_{\mathbb{x}}}U_{l}\in\mathbb{Z}_{N}$. Let $\varphi:\mathbb{Z}_{N}\to\mathbb{R}^{2}$ be the map that identifies $\mathbb{Z}_{N}$ with the complex roots of unity, and then regards them as points in $\mathbb{R}^{2}$. The smeared Polyakov loop for a given configuration is then given by taking the spatial average and interpreting it in $\mathbb{R}^{2}$:
$\overline{P}=\frac{1}{L^{3}}\sum_{\mathbb{x}}\varphi(P_{\mathbb{x}})\in\mathbb{R}^{2}.$
One immediate consequence of this definition is that the smeared Polyakov loop is contained in the $N$-gon spanned by the $N$-th roots of unity.
$e^{i\frac{2\pi}{5}}$$e^{i\frac{4\pi}{5}}$$e^{-i\frac{2\pi}{5}}$$e^{-i\frac{4\pi}{5}}$11
The histogram analysis was then reduced to constructing a density scatter plot of the measurement values within the $N$-gon, with each configuration contributing a single point.
The smeared Polyakov loop presents an $N$-fold rotational symmetry, corresponding to the $\mathbb{Z}_{N}$ 1-form symmetry introduced at the beginning. This is best illustrated in 2D. Let $H$ be an N1N-1 dimensional hyperplane on the dual lattice perpendicular to the time direction. In 2D, $H$ reduces to a line. The generator of the $\mathbb{Z}_{N}$ 1-form symmetry then acts on the lattice by shifting each intersecting link by 1 in $\mathbb{Z}_{N}$. As $H$ is assumed to be orthogonal to the time direction, this implies that every temporal Polyakov loop has a link crossing $H$, and thus that $P_{\mathbb{x}}\to P_{\mathbb{x}}+1\bmod N$ under the generator.